John is a door to door vacuum salesman. His weekly salary, s, is 2,000 pesos plus 50 pesos for each vacuum he sells. This can be written as S=2,000+50v, where v is the number of vacuum sold. If john earns 5000 for a weeks work, how many vacuums did he sell?
1
Expert's answer
2021-09-12T23:52:17-0400
"5000=2000+50v"

"50v=5000-2000"

"\dfrac{50v}{50}=\dfrac{3000}{50}"

"v=60"

John sold 50 vacuums for a weeks work.